# Towards Robust Recommendation: A Review and an Adversarial Robustness Evaluation Library

### v1

@@ Abstract
1 Introduction
With the explosive growth of information and data on the Internet, recommender systems have become a crucial tool for addressing the problem of information overload. These systems excel at solving the problem of matching users with relevant items, making them widely applicable across various domains, including online shopping platforms, video content recommendations, social networks, news portals, job recruitment platforms, and even in financial and healthcare systems.
While recommender systems offer a win-win situation for both users and businesses, it is evident that recommender systems are vulnerable to manipulation through malicious attacks. Kaya et al. (2023) [1] evaluated the robustness of three item-based multi-criteria recommender systems using the “PIA” attack model. The experimental results demonstrate that multicriteria top-n recommender systems are vulnerable to manipulation. Furthermore, He et al. (2018) [2] evaluated the performance of MF-BPR by introducing random noise and adversarial noise. The results indicate that both types of noise degrade the performance of MF-BPR, and adversarial noise has a greater impact. Furthermore, some studies [3, 4, 5] suggest that non-malicious factors such as sparse training data, sample imbalance, and unreasonable assumptions can also impact the performance of recommender systems. Therefore, the research on recommender systems has evolved beyond merely pursuing accuracy, with robustness becoming an important evaluation metric for recommender tasks.
Therefore, in this comprehensive review, we have extensively collected relevant works on recommender system robustness, categorized and introduced them, including common attack methods, malicious attack detection methods, robustness algorithms, and some non-adversarial robustness methods. Additionally, we have presented commonly used datasets and evaluation metrics in the task of recommender system robustness. Furthermore, we have implemented an adversarial robustness evaluation library, ShillingREC. ShillingREC integrates common attack models and recommendation models, enabling fair and efficient evaluation of these models.
1.1 Main Contributions
@@
1.2 Difference with Existing Surveys
@@
Figure 1: Recommender system Robustness Taxonomy Diagram
1.3 Structure of the Survey
Section 2 of this review presents the taxonomy for the robust recommender systems. Section 3 focuses on the adversarial robustness of recommender systems, specifically addressing shilling attacks. In Section 3.1, we introduce the concept and general definition of shilling attacks in the context of recommender systems. Section 3.2 categorizes shilling attack methods from different perspectives, followed by detailed discussions of specific methods in subsequent chapters. Section 4 delves into defense strategies against adversarial attacks on recommender systems, classifying and introducing various defense methods. Section 5 addresses the non-adversarial robustness of recommender systems. In Section 5.1, we define non-adversarial robustness and discuss its underlying causes, followed by a presentation of methods to enhance the non-adversarial robustness of recommender systems in subsequent chapters. Section 6 introduces commonly used datasets for evaluating the robustness of recommender systems. Section 7 elaborates on the evaluation metrics commonly employed in assessing the robustness of recommender systems. Section 8 introduces the ShillingREC library for evaluating shilling attacks and defense mechanisms in recommender systems. Section 9 outlines the current challenges and future research directions in the robustness of recommender systems. Finally, a concluding section summarizes the key points discussed throughout the review.
2 Taxonomy for Robust Recommendation
@@
2.1 Adversarial robustness
The emergence of robustness challenges in recommender systems can be attributed to both model properties and external factors.
@@ 2.1 Adversarial robustness
2.2 Non-adversarial robustness
In addition to malicious attacks targeting recommender systems, several non-malicious factors can also affect their performance, such as data sparsity, natural noise in training data, and data imbalance. The ability of recommender systems to withstand such negative effects is referred to as their non-adversarial robustness. A detailed discussion of the non-adversarial robustness of recommender systems is provided in Section 5.
3 Attack in adversarial robustness
3.1 Concept and General definition
@@ 3.2.3 Attack Method
TABLE I: Statistics of shilling attack approaches.
Category Method Data type RS task Rank strategy Metric Surrogate RS Attack knowledge Time
Target data Target recommend mode Target model parameters
@@
3.3 Heuristic-based attack
Heuristic attack methods employ predefined fixed strategies and statistical features to generate fake configuration profiles, which are then injected into recommender systems to achieve adversarial attacks. Such attack methods are simple to use, requiring minimal computational resources and time. They are also versatile, capable of targeting various recommendation models and scenarios. Common heuristic attack methods include random attack, average attack, bandwagon attack, love/hate attack, and segment attack.
In the random attack [18], the attacker initially needs to obtain the mean and standard deviation of all ratings in the dataset to construct a normal distribution for the rating scores. Subsequently, for each fake user, the attacker randomly selects n filler items based on hyperparameter settings. Ratings for these items are then assigned by sampling from the distribution of the rating data and discretizing them into specific scores. The rating of the target item is adjusted according to the specific attack objective, either to maximize or minimize it.
In the average attack [18], the attacker is required to calculate the mean and standard deviation of ratings for each item in the dataset and construct a normal distribution for its ratings. Similar to the Random Attack, the Average Attack then selects n filler items, and for each specific item, assigns values by sampling from the distribution of ratings for that item and discretizing them into specific scores.
The Bandwagon Attack, also known as the Popular Attack [19], incorporates a mixture of popular items (those with higher average ratings) and regular items in selecting filler items based on the specific scenario. Subsequently, it calculates the normal distribution of all ratings in the dataset and assigns values in a manner similar to the random attack setting.
In the Love/hate Attack, specifically taking the example of the push attack [19], for each fake user, the attacker randomly selects n filler items based on hyperparameter settings. Subsequently, the attacker sets the ratings for filler items to the minimum score and the target item’s rating to the maximum score.
The Segment Attack, as described in [20], aims to enhance promotion effectiveness by targeting users with similar tastes. Specifically, Segment Attack injects biased fake profiles into the recommender system, thereby constructing a set of fake data with high ratings for specific items that are similar in preference to the target items. Meanwhile, low ratings are assigned to other items. While heuristic methods are simple, fast, and widely applicable, their implementation principles suggest that they may lack precision, leading to suboptimal attack effectiveness. Moreover, the use of a singular generation strategy in heuristic attacks can make them prone to detection, reducing their stealthiness.
3.4 Optimization based attack
@@
3.4.1 Single-level optimization
@@
Therefore, to address this optimization problem, the Single-level optimization method requires an understanding of the structure and parameters of the victim model.
@@
3.4.2 Bi-level optimization
@@ 3.5 GAN based attack
In the domain of adversarial attacks on recommender systems, optimization-based adversarial attack methods often focus excessively on the harmfulness of the attack, neglecting the importance of concealing the attack. To achieve a covert and widely applicable adversarial attack, some studies have introduced GAN technology into the field of adversarial attacks. The DCGAN[25] method, proposed by Christakopoulou and Banerjee, is the first to utilize GAN technology for adversarial attacks. Specifically, the DCGAN generator takes the random noise input into convolutional units and produces fake profiles that conform to the distribution of real users. The discriminator is employed to identify the authenticity of the generated data. Additionally, to address the challenge of directly obtaining gradients for optimising attack objectives, zero-order optimization methods are used to estimate the gradients of the attack target, ultimately achieving a highly effective covert adversarial attack.
@@ 3.6 RL-based attack
4 Defense in adversarial robustness
@@
Lastly, leveraging various robustness algorithms, it is possible to develop robust recommendation models. These models are designed to minimize noise and learn genuine user representations during the learning process. By incorporating robustness measures into the model’s architecture, it becomes more resilient to shilling attacks and other adversarial influences.The schematic diagram illustrating the robustness defense mechanism is depicted in Figure 5.
@@ 4.1.1 Supervised methods
For instance, Zhou et al.[31] proposed an SVM-based shilling attack detection method called SVM-TIA. This method trains a supervised SVM classifier and employs the Borderline-SMOTE method to balance the dataset while using target-item analysis to reduce false positive rates. Tong et al.[32] introduced a CNN-based shilling attack detector for social-aware network collaborative recommender systems, known as CNN-SAD. CNN-SAD consists of a convolutional layer and a pooling layer, using sigmoid as the activation function. This represents the first application of deep learning techniques to detect recommendation attacks. Zhou[33] proposed a classifier called DL-DRA that eliminates the need for handcrafted features. DL-DRA first transforms rating vectors into rating matrices for each user. It then employs bicubic interpolation to adjust the size of rating matrices, reducing sparsity. Finally, a CNN network is constructed to detect fake users directly from this processed data. Ebrahimian et al.[34] introduced a mixed model that uses both CNN and RNN to detect shilling attacks, including CNN-LSTM and CNN-GRU. These methods can directly extract features from rating data and model the temporal and spatial information in the recommender system ratings, achieving end-to-end detection capabilities.
4.1.2 Unsupervised methods
@@
4.1.3 Semi-supervised methods
@@ 4.2 Robust algorithms
In light of these limitations, robustness algorithms for recommender systems have garnered significant research attention. Robustness algorithms aim to learn user representations that capture genuine preferences even in the presence of noisy data. Based on my research, this section categorizes mainstream recommender system robustness methods into four classes: adversarial training-based methods, contrastive learning-based methods, trust-aware recommendation methods, and model improvement-based methods.
4.2.1 Model based
@@
Due to MF techniques being one of the most successful and widely applied technologies in recommender systems, a multitude of robustness algorithms have emerged for MF.
@@
For instance, Cheng and Hurley (2010)introduced LTSMF[41] , which utilizes least squares for MF. Experimental results indicate that LTSMF exhibits stronger robustness compared to MF methods based on M-estimation. Yi et al. (2014) proposed the RCR-KTM method[42], which calculates user suspicion through neighborhood recommendations and combines it with MF based on Tukey’s M-estimator for recommendation. Yu et al. (2017) presented a robust recommendation method called RRA-KMF[43], which is based on kernel MF. RRA-KMF constructs a robust kernel MF model without square error terms using the rating matrix and kernel distances. Fadaee et al. (2018) introduced Chiron[44], a recommendation method that combines neighborhood-based and MF approaches. Chiron leverages both neighborhood information and MF for robust recommendations. Wu et al. (2020) proposed SL1-LF[45], which replaces the L2 norm in MF with a smooth L1 norm latent factor model. Experimental results demonstrate that the SL1-LF model exhibits robustness against anomalous data. These robustness algorithms tailored to MF contribute significantly to enhancing the robustness of recommender systems.
Recent research has shown that recommender systems based on Graph Neural Networks (GNNs) are susceptible to noise interference, which can impact their recommendation performance. Contrastive learning methods have emerged as effective approaches for enhancing the robustness of GNN-based recommender systems by maximizing the consistency of different views of the same node and learning users’ true preferences.
@@
For instance, Wu et al. (2021) proposed Self-Enhanced Graph Learning (SGL)[46], which supplements classical recommendation supervision tasks with auxiliary self-supervised tasks to strengthen node representation learning. Specifically, SGL generates multiple views of a node and maximizes the consistency between different views of the same node through self-supervised learning. Yang et al. (2022)[4] leveraged additional signals from knowledge graphs to supervise cross-graph contrastive learning. This approach improved bias reduction for user-item pairs during gradient descent and mitigated noise. It demonstrated good performance in scenarios involving sparse interactions, long-tail data, and noise. Zhu et al. (2023)[47] introduced Knowledge-Refined Denoising Network (KRDN), which combines adaptive knowledge refinement and contrastive denoising while removing irrelevant relationships from knowledge graphs and noise interactions from implicit feedback. In the knowledge refinement part, an adaptive pruning strategy is employed to extract high-quality triplets for additional knowledge in recommendations. In the contrastive denoising part, noise is identified based on the difference between collaborative signals and knowledge signals to learn users’ true preferences. Fan et al. (2023)[48] improved contrastive learning losses by applying Wasserstein distance measurements to quantify mutual information between enhanced sequences, thereby enhancing the robustness of sequence recommendations.
@@
In addition to the aforementioned approaches, the advent of denoising autoencoders has provided a new avenue for robustness algorithms in recommender systems. Wu et al. (2016) introduced the concept of using Collaborative Denoising Autoencoders (CDAE)[49] to mitigate observed random noise in interactions. They employed several linear layers to reconstruct the original dataset, thereby enhancing the model’s ability to resist noise.
4.2.2 Adversarial training
@@
4.2.3 Trust-aware recommender
In addition to model-based methods, trust information is also widely employed in robustness algorithms for recommender systems. These approaches aim to identify trusted users or trusted items for recommending to users, thereby reducing the impact of malicious attacks.

### v3

@@ Abstract
I Introduction
With the explosive growth of information and data on the Internet, recommender systems have become a crucial tool for addressing the problem of information overload. These systems excel at solving the problem of matching users with relevant items, making them widely applicable across various domains, including online shopping platforms, video content recommendations, social networks, news portals, and even in financial and healthcare systems.
While recommender systems offer a win-win situation for both users and businesses, it is evident that recommender systems are vulnerable to manipulation through malicious attacks. Kaya et al. (2023) [1] evaluated the robustness of three item-based multi-criteria recommender systems using the “PIA” attack model. The experimental results demonstrate that multicriteria top-n recommender systems are vulnerable to manipulation. In real-world scenarios, Amazon experienced a massive data breach involving fake reviews, which exposed more than 200,000 users and 13 million records, revealing widespread review manipulation on the platform [^1] https://www.techcentral.ie/data-breach-exposes-widespread-fake-reviews-on-amazon. Such attacks not only mislead consumers and damage platform credibility but also result in substantial economic losses. According to a report by CHEQ, fake reviews cause nearly \$152 billion in global economic damage annually [^2] https://www.saleswarp.com/fake-reviews-and-what-amazon-is-doing-about-it/.
Furthermore, some studies [2, 3, 4] suggest that non-malicious factors such as sparse training data, sample imbalance, and unreasonable assumptions can also impact the performance of recommender systems. Therefore, the research on recommender systems has evolved beyond merely pursuing accuracy, with robustness becoming an important evaluation metric for recommender tasks.
Therefore, in this comprehensive review, we have extensively collected relevant works on recommender system robustness, categorized and introduced them, including common attack methods, malicious attack detection methods, robust algorithms, and some non-adversarial robustness methods. Additionally, we have presented commonly used datasets and evaluation metrics in the task of recommender system robustness. Furthermore, we have implemented an adversarial robustness evaluation library, ShillingREC. ShillingREC integrates common attack models and recommendation models, enabling fair and efficient evaluation of these models.
I-A Main Contributions
@@
I-B Difference with Existing Surveys
@@
Figure 1: Recommender system robustness taxonomy diagram.
I-C Structure of the Survey
Section II of this review presents the taxonomy for the robust recommender systems. Section III focuses on the adversarial robustness of recommender systems, specifically addressing shilling attacks. In Section III-A, we introduce the concept and general definition of shilling attacks in the context of recommender systems. Section III-B categorizes shilling attack methods from different perspectives, followed by detailed discussions of specific methods in subsequent chapters. Section IV delves into defense strategies against adversarial attacks on recommender systems, classifying and introducing various defense methods. Section V addresses the non-adversarial robustness of recommender systems. In Section V-A, we define non-adversarial robustness and discuss its underlying causes, followed by a presentation of methods to enhance the non-adversarial robustness of recommender systems in subsequent chapters. Section VI introduces commonly used datasets for evaluating the robustness of recommender systems. Section VII elaborates on the evaluation metrics commonly employed in assessing the robustness of recommender systems. Section VIII introduces the ShillingREC library for evaluating shilling attacks and defense mechanisms in recommender systems. Section IX outlines the current challenges and future research directions in the robustness of recommender systems. Finally, a concluding section summarizes the key points discussed throughout the review.
II Taxonomy for Robust Recommendation
@@
II-A Adversarial Robustness
The emergence of robustness challenges in recommender systems stems from both model properties and external factors.
@@ II-A Adversarial Robustness
II-B Non-adversarial Robustness
In addition to malicious attacks targeting recommender systems, several non-malicious factors can also affect their performance, such as data sparsity, natural noise in training data, and data imbalance. The ability of recommender systems to withstand such negative effects is referred to as their non-adversarial robustness. A detailed discussion of the non-adversarial robustness of recommender systems is provided in Section V.
III Attack in Adversarial Robustness
III-A Concept and General Definition
@@ III-B3 Attack Method
TABLE I: Statistics of Shilling Attack Approaches
Category Method Data type RS task Rank strategy Metric Surrogate model Attack knowledge Time
Target
recommend
model
Target
model
parameters
@@
III-C Heuristic-based Attack
Heuristic-based attacks generate fake profiles using predefined fixed strategies and statistical features. These methods are straightforward, resource-efficient, and broadly applicable, allowing them to target various recommendation models and contexts. Common heuristic-based attack techniques include random attacks, average attacks, bandwagon attacks, love/hate attacks, and segment attacks.
In the random attack strategy [17], the first step is to assign the highest (push) or lowest (nuke) rating to the target item for all fake users. Then, the global rating mean and standard deviation are computed to establish a normal distribution for the rating scores. Finally, each fake user randomly selects n filler items, and the ratings for these items are randomly sampled from the constructed normal distribution.
In the average attack [17], the first step is to assign the highest (push) or lowest (nuke) rating to the target item for all fake users. Next, the mean and standard deviation of the ratings are calculated for each item to establish a normal distribution for its rating scores. Finally, each fake user randomly selects n filler items, and ratings are assigned by sampling from the normal distribution of the corresponding item’s rating scores.
Bandwagon Attack [18], or Popular Attack, resembles random attacks. It assigns maximum or minimum ratings to target items for all fake users and calculates the global mean and standard deviation to create a normal distribution of rating scores. The key distinction is in filler item selection: the Bandwagon Attack mixes a proportion of popular items (with higher average ratings) with random items, sampling rating scores from the established normal distribution.
The Love/Hate Attack [18] operates without prior knowledge of the data. In a push attack, all fake users are assigned the maximum rating for the target item, while each fake user randomly selects n filler items and assigns them the minimum rating. Although Love/Hate Attack is effective against user-based collaborative filtering, its straightforward filler strategy makes it relatively easy to detect.
Segment Attack [19] is a covert attack method designed to effectively recommend a target item to interested users. It assigns the highest rating to the target item for all fake users and selects several similar items as the selected items, also giving them the highest ratings. In addition, n filler items are randomly chosen and assigned the lowest ratings. Although Segment Attack requires more prior knowledge than other methods, its rating patterns closely resemble those of legitimate users, enhancing its subtlety.
While heuristic methods are simple, fast, and widely applicable, their implementation principles suggest that they may lack precision, leading to suboptimal attack effectiveness. Moreover, the use of a singular generation strategy in heuristic attacks can make them prone to detection, reducing their stealthiness.
III-D Optimization-based Attack
@@
III-D1 Single-level Optimization
@@
Therefore, to address this optimization problem, the Single-level optimization method generally requires an understanding of the structure and parameters of the victim model.
@@
Inspired by SRWA, Huang et al. (2021) [21] proposed a poisoning attack method for deep learning-based recommender systems called DL Attack. DL Attack utilizes a loss function to approximate the hit ratio and constructs a poison model to simulate a compromised deep learning recommender system. Ultimately, it builds filler items and their ratings for fake users through a two-stage training process of the poison model, employing a one-by-one approach.
III-D2 Bi-level Optimization
@@ III-E GAN-based Attack
In the domain of adversarial attacks on recommender systems, optimization-based adversarial attack methods often focus excessively on the harmfulness of the attack, neglecting the importance of concealing the attack. To achieve a covert and widely applicable adversarial attack, some studies have introduced GAN technology [25] into the field of adversarial attacks. The DCGAN[26] method, proposed by Christakopoulou and Banerjee, is the first to utilize GAN technology for adversarial attacks. Specifically, the DCGAN generator takes the random noise input into convolutional units and produces fake profiles that conform to the distribution of real users. The discriminator is employed to identify the authenticity of the generated data. Additionally, to address the challenge of directly obtaining gradients for optimising attack objectives, zero-order optimization methods are used to estimate the gradients of the attack target, ultimately achieving a highly effective covert adversarial attack.
@@ III-F RL-based Attack
IV Defense in Adversarial Robustness
@@
Lastly, leveraging various robust algorithms, it is possible to develop robust recommendation models. These models are designed to minimize noise and learn genuine user representations during the learning process. By incorporating robustness measures into the model’s architecture, it becomes more resilient to shilling attacks and other adversarial influences.The schematic diagram illustrating the robustness defense mechanism is depicted in Figure 5.
@@ IV-A1 Supervised Methods
For instance, Zhou et al.[32] proposed an SVM-based shilling attack detection method called SVM-TIA. This method trains a supervised SVM classifier and employs the Borderline-SMOTE method to balance the dataset while using target-item analysis to reduce false positive rates. Tong et al.[33] introduced a CNN-based shilling attack detector for social-aware network collaborative recommender systems, known as CNN-SAD. CNN-SAD consists of a convolutional layer and a pooling layer, using sigmoid as the activation function. This represents the first application of deep learning techniques to detect recommendation attacks. Ebrahimian et al.[34] introduced a mixed model that uses both CNN and RNN to detect shilling attacks, including CNN-LSTM and CNN-GRU. These methods can directly extract features from rating data and model the temporal and spatial information in the recommender system ratings, achieving end-to-end detection capabilities.
IV-A2 Unsupervised Methods
@@
IV-A3 Semi-supervised Methods
@@ IV-B Robust Algorithms
In light of these limitations, robust algorithms for recommender systems have garnered significant research attention. Robust algorithms aim to learn user representations that capture genuine preferences even in the presence of noisy data. Based on my research, this section categorizes mainstream recommender system robustness methods into three classes: model-based methods, adversarial training-based methods, and trust-aware recommendation methods.
IV-B1 Model Based
@@
Due to MF techniques being one of the most successful and widely applied technologies in recommender systems, a multitude of robust algorithms have emerged for MF.
@@
For instance, Cheng and Hurley (2010)introduced LTSMF[41] , which utilizes least squares for MF. Experimental results indicate that LTSMF exhibits stronger robustness compared to MF methods based on M-estimation. Yi et al. (2014) proposed the RCR-KTM method[42], which calculates user suspicion through neighborhood recommendations and combines it with MF based on Tukey’s M-estimator for recommendation. Yu et al. (2017) presented a robust recommendation method called RRA-KMF[43], which is based on kernel MF. RRA-KMF constructs a robust kernel MF model without square error terms using the rating matrix and kernel distances. Fadaee et al. (2018) introduced Chiron[44], a recommendation method that combines neighborhood-based and MF approaches. Chiron leverages both neighborhood information and MF for robust recommendations. Wu et al. (2020) proposed SL1-LF[45], which replaces the L2 norm in MF with a smooth L1 norm latent factor model. Experimental results demonstrate that the SL1-LF model exhibits robustness against anomalous data. These robust algorithms tailored to MF contribute significantly to enhancing the robustness of recommender systems.
Some researchs [46, 47, 48] suggest that GNN-based recommender systems are vulnerable to malicious noise due to their recursive message passing schema. To address this issue, certain works have introduced strategies such as graph regularization and contrastive learning to enhance the robustness of GNN-based recommendation models.
@@
For example, Chen et al. (2021) proposed the Structured Graph Convolutional Network (SGCN) [46], which enhances robustness against incomplete and noisy user-item graphs. SGCN achieves this by introducing random binary masks at each layer to prune noisy edges and applying nuclear norm regularization to maintain the low-rank characteristics of the graph. Wu et al. (2021) proposed Self-Enhanced Graph Learning (SGL)[49], which supplements classical recommendation supervision tasks with auxiliary self-supervised tasks to strengthen node representation learning. Specifically, SGL generates multiple views of a node and maximizes consistency between different views of the same node through self-supervised learning. Yang et al. (2022)[3] leveraged additional signals from knowledge graphs to supervise cross-graph contrastive learning. This approach improved bias reduction for user-item pairs during gradient descent and mitigated noise. It demonstrated good performance in scenarios involving sparse interactions, long-tail data, and noise. Zhu et al. (2023)[50] introduced the Knowledge-Refined Denoising Network (KRDN), which combines adaptive knowledge refinement and contrastive denoising while removing irrelevant relationships from knowledge graphs and noise interactions from implicit feedback. In the knowledge refinement part, an adaptive pruning strategy is employed to extract high-quality triplets for additional knowledge in recommendations. In the contrastive denoising part, noise is identified based on the difference between collaborative signals and knowledge signals to learn users’ true preferences. Fan et al. (2023)[51] improved contrastive learning losses by applying Wasserstein distance measurements to quantify mutual information between enhanced sequences, thus enhancing the robustness of sequence recommendations.
@@
In addition to the aforementioned approaches, denoising autoencoders have opened a new avenue for robust recommender algorithms. Wu et al. (2016) introduced Collaborative Denoising Autoencoders (CDAE) [52] to mitigate random noise in observed interactions. They employed several linear layers to reconstruct the original data, enhancing the model’s resistance to noise.
IV-B2 Adversarial Training
@@
IV-B3 Trust-aware Recommender
In addition to model-based methods, trust information is also widely employed in robust algorithms for recommender systems. These approaches aim to identify trusted users or trusted items for recommending to users, thereby reducing the impact of malicious attacks.